# Navigation World Models

### v1

@@ Abstract
Navigation is a fundamental skill of agents with visual-motor capabilities. We introduce a Navigation World Model (NWM), a controllable video generation model that predicts future visual observations based on past observations and navigation actions. To capture complex environment dynamics, NWM employs a Conditional Diffusion Transformer (CDiT), trained on a diverse collection of egocentric videos of both human and robotic agents, and scaled up to 1 billion parameters. In familiar environments, NWM can plan navigation trajectories by simulating them and evaluating whether they achieve the desired goal. Unlike supervised navigation policies with fixed behavior, NWM can dynamically incorporate constraints during planning. Experiments demonstrate its effectiveness in planning trajectories from scratch or by ranking trajectories sampled from an external policy. Furthermore, NWM leverages its learned visual priors to imagine trajectories in unfamiliar environments from a single input image, making it a flexible and powerful tool for next-generation navigation systems.
[…]
[Project Page]https://amirbar.net/nwm
@@ 1 Introduction
When human agents plan, they often imagine their future trajectories considering constraints and counterfactuals. On the other hand, current state-of-the-art robotics navigation policies (Sridhar et al., 2024; Shah et al., 2023) are “hard-coded”, and after training, new constraints cannot be easily introduced (e.g, “no left turns”). Another limitation of current supervised visual navigation models is that they cannot dynamically allocate more computational resources to address hard problems. We aim to design a new model that can mitigate these issues.
@@
In this work, we propose a Navigation World Model (NWM), trained to predict the future representation of a video frame based on past frame representation(s) and action(s) (see Figure 1(a)). NWM is trained on video footage and navigation actions collected from various robotic agents. After training, NWM is used to plan novel navigation trajectories by simulating potential navigation plans and verifying if they reach a target goal (see Figure 1(b)). To evaluate its navigation skills, we test NWM in known environments, assessing its ability to plan novel trajectories either independently or by ranking an external navigation policy. In the planning setup, we use NWM in a Model Predictive Control (MPC) framework, optimizing the action sequence that enables NWM to reach a target goal. In the ranking setup, we assume access to an existing navigation policy, such as NoMaD (Sridhar et al., 2024), which allows us to sample trajectories, simulate them using NWM, and select the best ones. Our NWM achieves competitive standalone performance and state-of-the-art results when combined with existing methods.
@@
NWM is conceptually similar to recent diffusion-based world models for offline model-based reinforcement learning, such as DIAMOND (Alonso et al., ) and GameNGen (Valevski et al., 2024). However, unlike these models, NWM is trained across a wide range of environments and embodiments, leveraging the diversity of navigation data from robotic and human agents. This allows us to train a large diffusion transformer model capable of scaling effectively with model size and data to adapt to multiple environments. Our approach also shares similarities with Novel View Synthesis (NVS) methods like NeRF (Mildenhall et al., 2021), Zero-1-2-3 (Liu et al., 2023), and GDC (Van Hoorick et al., 2024), from which we draw inspiration. However, unlike NVS approaches, our goal is to train a single model for navigation across diverse environments and model temporal dynamics from natural videos, without relying on 3D priors.
@@
To learn a NWM, we propose a novel Conditional Diffusion Transformer (CDiT), trained to predict the next image state given past image states and actions as context. Unlike a DiT (Peebles and Xie, 2023), CDiT’s computational complexity is linear with respect to the number of context frames, and it scales favorably for models trained up to $1B$ parameters across diverse environments and embodiments, requiring 4×4$\times$4 × fewer FLOPs compared to a standard DiT while achieving better future prediction results.
@@
In unknown environments, our results show that NWM benefits from training on unlabeled, action- and reward-free video data from Ego4D. Qualitatively, we observe improved video prediction and generation performance on single images (see Figure 1(c)). Quantitatively, with additional unlabeled data, NWM produces more accurate predictions when evaluated on the held-out Stanford Go (Hirose et al., 2018) dataset.
@@ 2 Related Work
Goal conditioned visual navigation is an important task in robotics requiring both perception and planning skills (Sridhar et al., 2024; Shah et al., ; Pathak et al., 2018; Mirowski et al., 2022; Chaplot et al., ; Fu et al., 2022). Given context image(s) and an image specifying the navigation goals, goal-conditioned visual navigation models (Sridhar et al., 2024; Shah et al., ) aim to generate a viable path towards the goal if the environment is known, or to explore it otherwise. Recent visual navigation methods like NoMaD (Sridhar et al., 2024) train a diffusion policy via behavior cloning and temporal distance objective to follow goals in the conditional setting or to explore new environments in the unconditional setting. Previous approaches like Active Neural SLAM (Chaplot et al., ) used neural SLAM together with analytical planners to plan trajectories in the $3D$ environment, while other approaches like (Chen et al., ) learn policies via reinforcement learning. Here we show that world models can use exploratory data to plan or improve existing navigation policies.
@@
Differently than in learning a policy, the goal of a world model (Ha and Schmidhuber, 2018) is to simulate the environment, e.g, given the current state and action to predict the next state and an associated reward. Previous works have shown that jointly learning a policy and a world model can improve sample efficiency on Atari (Hafner et al., b, a; Alonso et al., ), simulated robotics environments (Seo et al., 2023), and even when applied to real world robots (Wu et al., 2023). More recently, Hansen et al. proposed to use a single world model that is shared across tasks by introducing action and task embeddings while Yang et al. ; Lin et al. (2024b) proposed to describe actions in language, and Bruce et al. (2024) proposed to learn latent actions. World models were also explored in the context of game simulation. DIAMOND (Alonso et al., ) and GameNGen (Valevski et al., 2024) propose to use diffusion models to learn game engines of computer games like Atari and Doom. Our work is inspired by these works, and we aim to learn a single general diffusion video transformer that can be shared across many environments and different embodiments for navigation.
@@
In computer vision, generating videos has been a long standing challenge (Kondratyuk et al., ; Blattmann et al., 2023; Girdhar et al., 2023; Yu et al., 2023; Ho et al., 2022; Tulyakov et al., 2018b; Bar-Tal et al., 2024). Most recently, there has been tremendous progress with text-to-video synthesis with methods like Sora (Brooks et al., 2024) and MovieGen (Polyak et al., 2024). Past works proposed to control video synthesis given structured action-object class categories (Tulyakov et al., 2018a) or Action Graphs (Bar et al., 2021). Video generation models were previously used in reinforcement learning as rewards (Escontrela et al., 2024), pretraining methods (Tomar et al., 2024), for simulating and planning manipulation actions (Finn and Levine, 2017; Liang et al., 2024) and for generating paths in indoor environments (Hirose et al., 2019b; Koh et al., 2021). Interestingly, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are useful both for video tasks like generation (Voleti et al., 2022) and prediction (Lin et al., 2024a), but also for view synthesis (Chan et al., 2023; Poole et al., ; Tung et al., 2025). Differently, we use a conditional diffusion transformer to simulate trajectories for planning without explicit $3$D representations or priors.
@@ 3.1 Formulation
Next, we turn to describe our NWM formulation. Intuitively, a NWM is a model that receives the current state of the world (e.g, an image observation) and a navigation action describing where to move and how to rotate. The model then produces the next state of the world with respect to the agent’s point of view.
We are given an egocentric video dataset together with agent navigation actions $D=\{(x_{0},a_{0},...,x_{T},a_{T})\}^{n}_{i=1}$, such that $x_{i}\in\mathbb{R}^{H\times W\times 3}$ is an image and $a_{i}=(u,\phi)$ is a navigation command given by translation parameter $u\in\mathbb{R}^{2}$ that controls forward/backward and right/left motion, as well as and a yaw rotation angle $\phi\in\mathbb{R}$ [^1] This can be naturally extended to three dimensions by having $u\in\mathbb{R}^{3}$ and $\theta\in\mathbb{R}^{3}$ defining yaw, pitch and roll. For simplicity we assume navigation on a flat surface with fixed pitch and roll.
The navigation actions ${a_{i}}$ can be fully observed (as in Habitat (Savva et al., 2019)), e.g, moving forward towards a wall will trigger a response from the environment based on physics, which will lead to the agent staying in place, whereas in other environments we compute the navigation action based on the change in the agent’s location.
@@
$s_{i}=enc_{\theta}(x_{i})$ $s_{\tau+1}\sim F_{\theta}(s_{\tau+1}\mid\mathbf{s_{\tau}},a_{\tau})$ (1)
Where $\mathbf{s_{\tau}}=({s_{\tau},...,s_{\tau-m}})$ are the past $m$ visual observations encoded via a pretrained VAE (Blattmann et al., 2023). Using a VAE has the benefit of working with compressed latents, allowing to decode predictions back to pixel space for visualization.
Due to the simplicity of this formulation, it can be naturally shared across environments and easily extended to more complex action spaces, like controlling a robotic arm. Different than Hafner et al. (a), we aim to train a single world model across environments and embodiments, without using task or action embeddings like in Hansen et al. .
The formulation in Equation 1 models action but does not allow control over the temporal dynamics. We extend this formulation with a time shift input $k\in[{T_{\text{min}}},{T_{\text{max}}}]$, setting $a_{\tau}=(u,\phi,k)$, thus now $a_{\tau}$ specifies the time change $k$, used to determine how many steps should the model move into the future (or past). Hence, given a current state $s_{\tau}$, we can randomly choose $k$, tokenize the corresponding video frame then $s_{\tau+1}$ to be the corresponding set of tokens. The navigation actions can then be approximated to be a summation from time $\tau$ to $\tau+k$:
$u_{\tau\rightarrow\tau+k}=\sum_{t=\tau}^{\tau+k}u_{t}$ $\phi_{\tau\rightarrow\tau+k}=\sum_{t=\tau}^{\tau+k}\phi_{t}\mod 2\pi$ (2)
@@ 3.2 Diffusion Transformer as World Model
CDiT enables time efficient autoregressive modeling by constraining the attention in the first attention block only to tokens from the target frame which is being denoised. To condition on tokens from past frames, we incorporate a cross attention layer, such that every query token from the current target attends to tokens from past frames, which are used as keys and values. The cross attention then contextualizes the representations using a skip connection layer.
To condition on continuous actions like the translation $u$, rotation $\phi$, and time shift $k$, along with the diffusion timestep $t$, we map each scaler to sine-cosine features, then apply a $2$-layer MLP, $G$, mapping it from $\mathbb{R}$ to $\mathbb{R}^{d}$. Finally we sum all embeddings into a single vector used for conditioning:
$\xi=G_{u}(\psi(u))+G_{\theta}(\psi(\theta))+G_{k}(\psi(k))+G_{t}(\psi(t))$ (3)
$\xi$ is then fed to an AdaLN (Xu et al., 2019) block to generate scale and shift coefficients that modulate the Layer Normalization (Lei Ba et al., 2016) outputs , as well as the outputs of the attention layers. To train on unlabeled data, we simple omit explicit navigation actions when computing $\xi$ (see Eq 3).
An alternative approach is to simply use DiT (Peebles and Xie, 2023), however, applying a DiT on the full input is computationally expensive. Denote $n$ the number of input tokens per frame, and $m$ the number of frames, and $d$ the token dimension. Scaled Multi-head Attention Layer (Vaswani, 2017) complexity is dominated by the attention term $O(m^{2}\cdot n^{2}\cdot d)$, which is quadratic with context length. In contrast, our CDiT block is dominated by the cross-attention layer complexity $O(m\cdot n^{2}\cdot d)$, which is linear with respect to the context, allowing us to use longer context size. We analyze these two design choices in Section 4. CDiT resembles the original Transformer Block (Vaswani, 2017), without applying expensive self-attention over the context tokens.
@@
Diffusion Training. In the forward process, noise is added to the target state $s_{\tau+1}$ according to a randomly chosen timestep $t\in\{1,\dots,T\}$. The noisy state $s_{\tau+1}^{(t)}$ can be defined as: $s_{\tau+1}^{(t)}=\sqrt{\alpha_{t}}s_{\tau+1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\{\alpha_{t}\}$ is a noise schedule controlling the variance. As $t$ increases, $s_{\tau+1}^{(t)}$ converges to pure noise. The backward process attempts to recover the original state representation $s_{\tau+1}$ from the noisy version $s_{\tau+1}^{(t)}$, conditioned on the context $\mathbf{s}_{\tau}$, the current action $a_{\tau}$, and the diffusion timestep $t$. We define $F_{\theta}(s_{\tau+1}|\mathbf{s}_{\tau},a_{\tau},t)$ as the neural network model parameterized by $\theta$ that predicts the noise added at each step. We follow the same noise schedule and hyperparameters as in DiT (Peebles and Xie, 2023).
Training Objective. The model is trained to minimize the mean-squared error between the noisy target $s_{\tau+1}^{(t)}$ and the predicted target $F_{\theta}(s_{\tau+1}^{(t)}|\mathbf{s}_{\tau},a_{\tau},t)$, aiming to learn the denoising process:
$\mathcal{L}_{\theta}=\mathbb{E}_{s_{\tau+1},a_{\tau},\mathbf{s}_{\tau},\epsilon,t}\left[\|s_{\tau+1}-F_{\theta}(s_{\tau+1}^{(t)}|\mathbf{s}_{\tau},a_{\tau},t)\|^{2}\right].$
In this objective, $t$ is sampled randomly, ensuring that the model learns to denoise frames at all levels of corruption. By minimizing this loss, the model learns to reconstruct $s_{\tau+1}$ from $s_{\tau+1}^{(t)}$ based on the context $\mathbf{s}_{\tau}$ and the action $a_{\tau}$, thus generating realistic future frames in a sequence.

### v2

@@ Abstract
Navigation is a fundamental skill of agents with visual-motor capabilities. We introduce a Navigation World Model (NWM), a controllable video generation model that predicts future visual observations based on past observations and navigation actions. To capture complex environment dynamics, NWM employs a Conditional Diffusion Transformer (CDiT), trained on a diverse collection of egocentric videos of both human and robotic agents, and scaled up to 1 billion parameters. In familiar environments, NWM can plan navigation trajectories by simulating them and evaluating whether they achieve the desired goal. Unlike supervised navigation policies with fixed behavior, NWM can dynamically incorporate constraints during planning. Experiments demonstrate its effectiveness in planning trajectories from scratch or by ranking trajectories sampled from an external policy. Furthermore, NWM leverages its learned visual priors to imagine trajectories in unfamiliar environments from a single input image, making it a flexible and powerful tool for next-generation navigation systems [^1] Project page: https://amirbar.net/nwm.
{strip}
@@ 1 Introduction
When human agents plan, they often imagine their future trajectories considering constraints and counterfactuals. On the other hand, current state-of-the-art robotics navigation policies [55, 53] are “hard-coded”, and after training, new constraints cannot be easily introduced (e.g. “no left turns”). Another limitation of current supervised visual navigation models is that they cannot dynamically allocate more computational resources to address hard problems. We aim to design a new model that can mitigate these issues.
@@
In this work, we propose a Navigation World Model (NWM), trained to predict the future representation of a video frame based on past frame representation(s) and action(s) (see Figure 1(a)). NWM is trained on video footage and navigation actions collected from various robotic agents. After training, NWM is used to plan novel navigation trajectories by simulating potential navigation plans and verifying if they reach a target goal (see Figure 1(b)). To evaluate its navigation skills, we test NWM in known environments, assessing its ability to plan novel trajectories either independently or by ranking an external navigation policy. In the planning setup, we use NWM in a Model Predictive Control (MPC) framework, optimizing the action sequence that enables NWM to reach a target goal. In the ranking setup, we assume access to an existing navigation policy, such as NoMaD [55], which allows us to sample trajectories, simulate them using NWM, and select the best ones. Our NWM achieves state-of-the-art standalone performance and competitive results when combined with existing methods.
@@
NWM is conceptually similar to recent diffusion-based world models for offline model-based reinforcement learning, such as DIAMOND [1] and GameNGen [66]. However, unlike these models, NWM is trained across a wide range of environments and embodiments, leveraging the diversity of navigation data from robotic and human agents. This allows us to train a large diffusion transformer model capable of scaling effectively with model size and data to adapt to multiple environments. Our approach also shares similarities with Novel View Synthesis (NVS) methods like NeRF [40], Zero-1-2-3 [38], and GDC [67], from which we draw inspiration. However, unlike NVS approaches, our goal is to train a single model for navigation across diverse environments and model temporal dynamics from natural videos, without relying on 3D priors.
@@
To learn a NWM, we propose a novel Conditional Diffusion Transformer (CDiT), trained to predict the next image state given past image states and actions as context. Unlike a DiT [44], CDiT’s computational complexity is linear with respect to the number of context frames, and it scales favorably for models trained up to $1B$ parameters across diverse environments and embodiments, requiring 4×4$\times$4 × fewer FLOPs compared to a standard DiT while achieving better future prediction results.
@@
In unknown environments, our results show that NWM benefits from training on unlabeled, action- and reward-free video data from Ego4D. Qualitatively, we observe improved video prediction and generation performance on single images (see Figure 1(c)). Quantitatively, with additional unlabeled data, NWM produces more accurate predictions when evaluated on the held-out Stanford Go [24] dataset.
@@ 2 Related Work
Goal conditioned visual navigation is an important task in robotics requiring both perception and planning skills [55, 51, 43, 41, 8, 15, 13]. Given context image(s) and an image specifying the navigation goals, goal-conditioned visual navigation models [55, 51] aim to generate a viable path towards the goal if the environment is known, or to explore it otherwise. Recent visual navigation methods like NoMaD [55] train a diffusion policy via behavior cloning and temporal distance objective to follow goals in the conditional setting or to explore new environments in the unconditional setting. Previous approaches like Active Neural SLAM [8] used neural SLAM together with analytical planners to plan trajectories in the $3D$ environment, while other approaches like [9] learn policies via reinforcement learning. Here we show that world models can use exploratory data to plan or improve existing navigation policies.
@@
Differently than in learning a policy, the goal of a world model [19] is to simulate the environment, e.g. given the current state and action to predict the next state and an associated reward. Previous works have shown that jointly learning a policy and a world model can improve sample efficiency on Atari [21, 20, 1], simulated robotics environments [50], and even when applied to real world robots [71]. More recently, [22] proposed to use a single world model that is shared across tasks by introducing action and task embeddings while [73, 37] proposed to describe actions in language, and [6] proposed to learn latent actions. World models were also explored in the context of game simulation. DIAMOND [1] and GameNGen [66] propose to use diffusion models to learn game engines of computer games like Atari and Doom. Our work is inspired by these works, and we aim to learn a single general diffusion video transformer that can be shared across many environments and different embodiments for navigation.
@@
In computer vision, generating videos has been a long standing challenge [32, 4, 17, 74, 29, 62, 3]. Most recently, there has been tremendous progress with text-to-video synthesis with methods like Sora [5] and MovieGen [45]. Past works proposed to control video synthesis given structured action-object class categories [61] or Action Graphs [2]. Video generation models were previously used in reinforcement learning as rewards [10], pretraining methods [59], for simulating and planning manipulation actions [11, 35] and for generating paths in indoor environments [26, 31]. Interestingly, diffusion models [54, 28] are useful both for video tasks like generation [69] and prediction [36], but also for view synthesis [7, 46, 63]. Differently, we use a conditional diffusion transformer to simulate trajectories for planning without explicit $3$D representations or priors.
@@ 3.1 Formulation
Next, we turn to describe our NWM formulation. Intuitively, a NWM is a model that receives the current state of the world (e.g. an image observation) and a navigation action describing where to move and how to rotate. The model then produces the next state of the world with respect to the agent’s point of view.
We are given an egocentric video dataset together with agent navigation actions $D=\{(x_{0},a_{0},...,x_{T},a_{T})\}^{n}_{i=1}$, such that $x_{i}\in\mathbb{R}^{H\times W\times 3}$ is an image and $a_{i}=(u,\phi)$ is a navigation command given by translation parameter $u\in\mathbb{R}^{2}$ that controls the change in forward/backward and right/left motion, as well as $\phi\in\mathbb{R}$ that controls the change in yaw rotation angle. [^2] This can be naturally extended to three dimensions by having $u\in\mathbb{R}^{3}$ and $\theta\in\mathbb{R}^{3}$ defining yaw, pitch and roll. For simplicity, we assume navigation on a flat surface with fixed pitch and roll.
The navigation actions ${a_{i}}$ can be fully observed (as in Habitat [49]), e.g. moving forward towards a wall will trigger a response from the environment based on physics, which will lead to the agent staying in place, whereas in other environments the navigation actions can be approximated based on the change in the agent’s location.
@@
$s_{i}=\text{enc}_{\theta}(x_{i})$ $s_{\tau+1}\sim F_{\theta}(s_{\tau+1}\mid\mathbf{s_{\tau}},a_{\tau})$ (1)
Where $\mathbf{s_{\tau}}=({s_{\tau},...,s_{\tau-m}})$ are the past $m$ visual observations encoded via a pretrained VAE [4]. Using a VAE has the benefit of working with compressed latents, allowing to decode predictions back to pixel space for visualization.
Due to the simplicity of this formulation, it can be naturally shared across environments and easily extended to more complex action spaces, like controlling a robotic arm. Different than [20], we aim to train a single world model across environments and embodiments, without using task or action embeddings like in [22].
The formulation in Equation 1 models action but does not allow control over the temporal dynamics. We extend this formulation with a time shift input $k\in[{T_{\text{min}}},{T_{\text{max}}}]$, setting $a_{\tau}=(u,\phi,k)$, thus now $a_{\tau}$ specifies the time change $k$, used to determine how many steps should the model move into the future (or past). Hence, given a current state $s_{\tau}$, we can randomly choose a timeshift $k$ and use the corresponding time shifted video frame as our next state $s_{\tau+1}$. The navigation actions can then be approximated to be a summation from time $\tau$ to $m=\tau+k-1$:
$u_{\tau\rightarrow m}=\sum_{t=\tau}^{m}u_{t}$ $\phi_{\tau\rightarrow m}=\sum_{t=\tau}^{m}\phi_{t}\mod 2\pi$ (2)
@@ 3.2 Diffusion Transformer as World Model
CDiT enables time-efficient autoregressive modeling by constraining the attention in the first attention block only to tokens from the target frame which is being denoised. To condition on tokens from past frames, we incorporate a cross-attention layer, such that every query token from the current target attends to tokens from past frames, which are used as keys and values. The cross-attention then contextualizes the representations using a skip connection layer.
To condition on the navigation action $a\in\mathbb{R}^{3}$, we first map each scalar to $\mathbb{R}^{\frac{d}{3}}$ by extracting sine-cosine features, then applying a $2$-layer MLP, and concatenating them into a single vector $\psi_{a}\in\mathbb{R}^{d}$. We follow a similar process to map the timeshift $k\in\mathbb{R}$ to $\psi_{k}\in\mathbb{R}^{d}$ and the diffusion timestep $t\in\mathbb{R}$ to $\psi_{k}\in\mathbb{R}^{d}$. Finally we sum all embeddings into a single vector used for conditioning:
$\xi=\psi_{a}+\psi_{k}+\psi_{t}$ (3)
$\xi$ is then fed to an AdaLN [72] block to generate scale and shift coefficients that modulate the Layer Normalization [34] outputs, as well as the outputs of the attention layers. To train on unlabeled data, we simply omit explicit navigation actions when computing $\xi$ (see Eq. 3).
An alternative approach is to simply use DiT [44], however, applying a DiT on the full input is computationally expensive. Denote $n$ the number of input tokens per frame, and $m$ the number of frames, and $d$ the token dimension. Scaled Multi-head Attention Layer [68] complexity is dominated by the attention term $O(m^{2}n^{2}d)$, which is quadratic with context length. In contrast, our CDiT block is dominated by the cross-attention layer complexity $O(mn^{2}d)$, which is linear with respect to the context, allowing us to use longer context size. We analyze these two design choices in Section 4. CDiT resembles the original Transformer Block [68], without applying expensive self-attention over the context tokens.
@@
Diffusion Training. In the forward process, noise is added to the target state $s_{\tau+1}$ according to a randomly chosen timestep $t\in\{1,\dots,T\}$. The noisy state $s_{\tau+1}^{(t)}$ can be defined as: $s_{\tau+1}^{(t)}=\sqrt{\alpha_{t}}s_{\tau+1}+\sqrt{1-\alpha_{t}}\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\{\alpha_{t}\}$ is a noise schedule controlling the variance. As $t$ increases, $s_{\tau+1}^{(t)}$ converges to pure noise. The reverse process attempts to recover the original state representation $s_{\tau+1}$ from the noisy version $s_{\tau+1}^{(t)}$, conditioned on the context $\mathbf{s}_{\tau}$, the current action $a_{\tau}$, and the diffusion timestep $t$. We define $F_{\theta}(s_{\tau+1}|\mathbf{s}_{\tau},a_{\tau},t)$ as the denoising neural network model parameterized by $\theta$. We follow the same noise schedule and hyperparams of DiT [44].
Training Objective. The model is trained to minimize the mean-squared between the clean and predicted target, aiming to learn the denoising process:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{s_{\tau+1},a_{\tau},\mathbf{s}_{\tau},\epsilon,t}\left[\|s_{\tau+1}-F_{\theta}(s_{\tau+1}^{(t)}|\mathbf{s}_{\tau},a_{\tau},t)\|_{2}^{2}\right].$
In this objective, the timestep $t$ is sampled randomly to ensure that the model learns to denoise frames across varying levels of corruption. By minimizing this loss, the model learns to reconstruct $s_{\tau+1}$ from its noisy version $s_{\tau+1}^{(t)}$, conditioned on the context $\mathbf{s}_{\tau}$ and action $a_{\tau}$, thereby enabling the generation of realistic future frames. Following [44], we also predict the covariance matrix of the noise and supervise it with the variational lower bound loss $\mathcal{L}_{\text{vlb}}$ [42].